OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. 0 = x2 + 5x + 6 0 = x2 + 4x + 4

OpenStudy (anonymous):

(x+3)(x+2)=0 (x+2)(x+2)=0 the solutions represent where the graph cuts the x-axis. the graph cuts the x-axis at 2 points. in the first equation, the graph cuts the x axis at at x=-3 and x=-2 in the second equation the graph cuts the x-axis at x=-2 (repeated root).